Suppose the risk-free rate is 1.21% and an analyst assumes a market risk premium of 5.44%. Firm A just paid a dividend of $1.23 per share. The analyst estimates the β of Firm A to be 1.45 and estimates the dividend growth rate to be 4.29% forever. Firm A has 257.00 million shares outstanding. Firm B just paid a dividend of $1.91 per share. The analyst estimates the β of Firm B to be 0.79 and believes that dividends will grow at 2.60% forever. Firm B has 194.00 million shares outstanding. What is the value of Firm A?
Answer #1

Cost of Equity of firm A =Risk Free rate+Beta*Market Risk Premium =1.21%+1.45*5.44% =9.098%
Price of stock A =Dividend Next Year/(Cost of Equity-Growth) =1.23*(1+4.29%)/(9.098%-4.29%) =26.6798460898502

The Value of firm A =Price of share*Number of shares =26.6798460898502*257,000,000=6,856,720,445.09